# TLUE: A Tibetan Language Understanding Evaluation Benchmark

### v3

@@ Appendix C Appendix: Extended Experiment
Privacy and Property PP
Response Rate RR
@@ C.1 Parameters Settings of LLMs
As shown in Table 5, regarding open-source model execution, the Qwen-2.5 Team (2024) and DeepSeek families DeepSeek-AI et al. (2024, 2025) were evaluated via their respective official APIs provided by the model developers. The LlaMA-3.1 Dubey et al. (2024) (8B, 70B, 405B) were accessed and run using the LlaMA-API platform. These details will also be clearly documented in the final version to enhance transparency and reproducibility.
@@ C.2 Experimental Details
Table 6 and Table 7 shows a significant performance drop in Ti-MMLU and Ti-SafetyBench, with most results falling below the random baseline, highlighting the difficulty of safety alignment in low-resource languages.
@@
LLM Version Method Avg. STEM Human Social Other China
@@
Model Version Method Avg. OFF UB PH MH IA EM PP
@@ Appendix H Appendix: Clarification on Model Comparison under Low Accuracy
Each prediction reflects whether a LLM can handle specific domains or linguistic constructions. To complement the quantitative findings in Section 5.6, we include a qualitative comparison in Table 15. This example contrasts two LLMs from the Qwen-2.5 series Team (2024) on the same Ti-MMLU question. Notably, the larger 72B model generates a fluent but incorrect explanation, while the 32B model selects the correct answer with a more concise yet accurate rationale. This case illustrates non-monotonic scaling behavior and differences in reasoning stability, highlighting the value of instance-level analysis even under low average performance.

### v5

@@ Appendix C Appendix: Extended Experiment
Privacy and Property $P$
Response Rate $R$
@@ C.1 Parameters Settings of LLMs
As shown in Table 5, regarding open-source model execution, the Qwen-2.5 Qwen Team (2024) and DeepSeek families DeepSeek-AI et al. (2024, 2025) were evaluated via their respective official APIs provided by the model developers. The LlaMA-3.1 Dubey et al. (2024) (8B, 70B, 405B) were accessed and run using the LlaMA-API platform. These details will also be clearly documented in the final version to enhance transparency and reproducibility.
@@ C.2 Experimental Details
Table 6 and Table 7 show a significant performance drop in Ti-MMLU and Ti-SafetyBench, with most results falling below the random baseline, highlighting the difficulty of safety alignment in low-resource languages.
@@
LLM Version
Method
Avg. STEM Human
Social
Other
China
@@
Model Version Method Avg. OFF UB PH MH IA EM $P$
@@ Appendix H Appendix: Clarification on Model Comparison under Low Accuracy
Each prediction reflects whether a LLM can handle specific domains or linguistic constructions. To complement the quantitative findings in Section 5.6, we include a qualitative comparison in Table 15. This example contrasts two LLMs from the Qwen-2.5 series Qwen Team (2024) on the same Ti-MMLU question. Notably, the larger 72B model generates a fluent but incorrect explanation, while the 32B model selects the correct answer with a more concise yet accurate rationale. This case illustrates non-monotonic scaling behavior and differences in reasoning stability, highlighting the value of instance-level analysis even under low average performance.